# Particle Acceleration in Collisionless Magnetically Arrested Disks

### v1

@@ Abstract
We present the first collisionless realization of black hole accretion consistent with a persistent magnetically arrested disk state. The accretion flow, consisting of an ion-electron disk plasma combined with magnetospheric pair creation effects, is simulated using first-principles general-relativistic particle-in-cell methods. The axisymmetric simulation is evolved over significant dynamical timescales during which a quasi-steady accretion state is reached with several magnetic flux eruption cycles. We include a realistic treatment of inverse Compton scattering and pair production, which allows for studying the interaction between the collisionless accretion flow and the pair-loaded jet. Our findings indicate that magnetic flux eruptions associated with equatorial magnetic reconnection within the black hole magnetosphere and the formation of spark gaps are locations of maximal particle acceleration. Flux eruptions, starting near the central black hole, can trigger Kelvin-Helmholtz-like vortices at the jet-disk interface that facilitate efficient mixing between disk and jet plasma. Transient periods of increased pair production following the magnetic flux eruptions and reconnection events are responsible for most of the highly accelerated particles.
@@
Supermassive black holes (BH) at the center of low-luminosity active galactic nuclei (AGN) accrete ion-lepton plasma whose high temperatures and low densities place it in the collisionless regime. Some of these BHs are known to launch jets [via, e.g., the Blandford-Znajek (BZ) mechanism; 1], but it is well-established that not all systems clearly display these features [2, 3, 4]. Examples can be found in the two low-luminosity AGN M87 and Sagittarius A, where M87 shows a prominent, large-scale jet and Sagittarius A only displays radiative characteristics typical for a compact jet [5, 6]. The Event Horizon Telescope has observed emission structures associated with millimeter-wavelength synchrotron radiation from near-horizon relativistic accretion flows for both systems [7, 8]. Although the ring-like features are likely produced by the bulk accretion flow, the question remains of how and where the high-energy particles and emission signatures are created.
@@
Many jetted black hole systems, including M87 and Sagittarius A, display $\gamma$-ray activity [9, 10, 11, 12, 13, 14, 15]. These systems can go through periods of increased (i.e., flaring) activity [16, 17, 18], which can be related to the dynamics of the central engine that drives jet launching mechanics and its associated plasma effects. An important source of the high-energy emission is related to the plasma dynamics within the BH magnetosphere, but it remains difficult to determine which mechanism is predominantly responsible and whether the collisionless accretion flow dynamics are influential. In this numerical study, we therefore aim to simultaneously assess the ability of accretion-driven and magnetospheric processes to accelerate particles to high, non-thermal energies.
@@
The Magnetically Arrested Disk [MAD; 19, 20] state is currently a preferred model for interpreting millimeter emission from the two aforementioned low-luminosity AGN, both with regard to the polarimetric signatures [21] and the origin of flares with associated orbital motion of hot spots[22, 23, 24, 25]. MAD episodes are characterized by flux eruptions, during which a limiting horizon-penetrating magnetic flux is reached and the accretion flow can be partially halted [e.g., 26, 27, 28, 29, 30, 31]. General-relativistic magnetohydrodynamical (GRMHD) models that bring about the MAD state are characterized by large-scale coherent magnetic field structures and significant matter reservoirs in the initialized accretion disks. However, GRMHD methods are technically not applicable in the collisionless regime and do not directly provide insight into the non-ideal processes associated with particle acceleration. It is therefore valuable to investigate these systems with methods that self-consistently take into account these processes while simultaneously building on the GRMHD findings.
@@
In this Letter, we employ general relativistic (radiative) particle-in-cell (GR[R]PIC) methods to undertake a global collisionless study combining initialized accretion flows and particle creation processes. We present a quasi-steady MAD accretion state and focus on interpreting the locations and mechanisms of particle acceleration within this environment. The system’s characteristics — highly magnetized, with large-scale coherent field structures — are reminiscent of the well-established MAD state in GRMHD models [e.g., 26, 27, 28, 29, 30, 31]. Previously, several kinetic studies of black hole magnetospheres were undertaken that mostly focused on pair-created plasma dynamics under magnetically dominated conditions [32, 33, 34, 35, 36, 37, 38] or spherically-symmetric Bondi-Hoyle accretion [39]. Here, we present the first realization of disk-like initial conditions in a collisionless scenario, consisting of a three-species (pair-ion) plasma with a significant ion-to-electron mass ratio. With the addition of a strong accretion flow component to the BH magnetosphere in the presented simulation, it becomes possible to better constrain flaring dynamics of AGN systems and gain insight into the edge-brightened mechanisms of AGN jets [40, 41, 42, 43, 44, 45].
@@
We use the GRRPIC code Zeltron [46, 32] to evolve an ion-electron accretion flow combined with a pair-filled black hole magnetosphere. The total simulation domain is denoted by $r\in[0.9r_{\mathrm{h}},27r_{\mathrm{g}}]$ and $\theta\in[0.01\pi,0.99\pi]$ with a resolution of $N_{r}\times N_{\theta}=1440^{2}$ for spherical Kerr-Schild coordinates that are logarithmically spaced in the $\hat{r}$-direction. Here, $r_{\mathrm{g}}=GM/c^{2}$ is the gravitational radius and $r_{\mathrm{h}}=(1+\sqrt{1-a_{\ast}^{2}})r_{\mathrm{g}}$ is the horizon radius, with gravitational constant $G$, black hole mass $M$, speed of light $c$, and dimensionless black hole spin $a_{\ast}=0.99$. The ion mass is set to $m_{i}=100\,m_{e}$, where $m_{e}$ is the electron mass. We initialize a uniform density ($8\,n_{\textrm{GJ}}$) wedge that spans $r\in[5,27]r_{\mathrm{g}}$, $\theta\in[\pi/3,2\pi/3]$, where $n_{\mathrm{GJ}}=B_{0}\Omega_{\mathrm{h}}/(4\pi ce)$ is the Goldreich-Julian density with $\Omega_{\mathrm{h}}=a_{\ast}/(r_{\mathrm{h}}^{2}+a_{\ast}^{2})$ [47]. In the inner region, i.e., $r\in[r_{\mathrm{h}},5\,r_{\mathrm{g}}]$, we initialize a uniform pair-plasma sphere with a density of $1\,n_{\mathrm{GJ}}$ to kickstart pair production near the hole. The magnetic field begins as a single large magnetic loop in the poloidal plane [see Supplemental Material; 48]; the ratio of thermal to magnetic pressure is $\beta\in[0.1,5]$ at initialization.
@@
The initial configuration of the ion-electron plasma is somewhat arbitrarily chosen and of little consequence to the formation of the relaxed, quasi-steady configuration that is achieved after the simulation is evolved for sufficient time [see Supplemental Material; 48]. No angular momentum is introduced for the accretion flow at initialization, which expedites the onset of a quasi-steady accretion state. For simplicity, we refer to the low angular-momentum accretion flow as the “disk”. The ion-electron disk plasma is replenished from a spherical shell corresponding to $r\in[25,27]r_{\mathrm{g}}$, $\theta\in[\pi/3,2\pi/3]$, where the injection occurs with a fixed rate when the density drops below the initialized disk density. The injection from the boundary emulates the inflow of matter from a large-scale accretion flow. Particles that pass through the outer boundary are removed, mimicking an open boundary [49].
@@
Electron-positron pairs are created via photon-photon annihilation of an inverse-Compton scattered $\gamma$-ray photons with a background photon field [48]. The inclusion of these particle creation effects provides insight into how and if the jet is loaded with pairs and whether these pairs are able to significantly interact at the jet-disk interface. We refer the reader to the Supplemental Material [48] (and references [50, 33]) for a more detailed description of the simulation specifics, such as the definition of the poloidal magnetic field initialization, pair production criteria, and characteristic scaling.
@@
Figure 2: Overview of the MAD accretion state for two moments in time. The displayed quantities are magnetization $\sigma_{\mathrm{h}}=B^{2}/\left[4\pi\sum_{s}\langle\gamma_{s}\rangle n_{s}m_{s}c^{2}\right]$, where $\langle\gamma_{s}\rangle$ is the cell-averaged Lorentz factor and $\rho_{s}=m_{s}\,n_{s}$ the mass density per species, $s\in\{i,e^{-},e^{+}\}$, normalized with the Goldreich-Julian mass density $\rho_{\mathrm{GJ}}=m_{e}\,n_{\mathrm{GJ}}$. Kelvin-Helmholtz vortices are seen to form along the jet-disk interface. The inset panels show the density-averaged energy per particle $\langle\xi\rangle=\sum_{s}\langle\gamma_{s}\rangle\rho_{s}c^{2}/\sum_{s}\rho_{s}$ for different regions of the simulation domain. The green contours outline the magnetic flux surfaces. The dashed blue contours denote the ergosphere boundary.
@@
The accretion flow dynamic is well summarized in the time series of horizon-penetrating fluxes, where we define the mass accretion rate $\dot{M}=-\iint_{A}\rho u^{r}\sqrt{-g}\mathrm{d}\theta\mathrm{d}\phi$ and magnetic flux $\Phi_{\mathrm{B}}=\tfrac{1}{2}\iint_{A}|B^{r}|\sqrt{h}\mathrm{d}\theta\mathrm{d}\phi$ on the horizon ($r=r_{\mathrm{h}}$). Here, $\rho$, $u^{\mu}$, $g$, $h$, and $B^{r}$ are the mass density, metric determinant, spatial metric determinant, and the radial component of magnetic field as seen by the local fiducial observers (FIDOs), respectively. The quantity known as the MAD parameter, defined as $\phi_{\mathrm{BH}}=\Phi_{\mathrm{B}}/\sqrt{\dot{M}}$, is known to saturate at ${\sim}50$ [51]. After it saturates, a flux eruption event occurs, where the magnetic flux is (partially) advected away from the BH horizon. This leads to the reconnection of the upper and lower hemispheres of the BH magnetosphere in the equatorial plane [29, 30, 31]. We also find that the pairs created by photon-photon annihilation leave an imprint on the system’s flux eruptions, as a significant (vacuum) gap can form during the peak of the flux eruption that is rapidly filled by a pair production cascade.
@@
In Fig. 1, we show direct evidence for a quasi-steady MAD accretion state achieved with GRRPIC methods. The top panel clearly outlines sharp decreases in the mass accretion rate $\dot{M}$, which are direct consequences of the flux eruption events, where the accretion flow is temporarily halted because of a magnetic field barrier created by the flux eruption. This relates to the saw-tooth structure seen for $\Phi_{\mathrm{B}}$, where the exponential decay is related to the reconnection event (or flux eruption). The general shape of growing $\Phi_{\mathrm{B}}$ with the saw-tooth structures of ${\sim}50\,r_{\mathrm{g}}/c$ in length combined with significant amplitude changes have not been seen before in other GRPIC simulations [cf. 34, 52, 39]. The MAD state, corresponding to reaching the limit $\phi_{\mathrm{BH}}\gtrsim 50$, is established from the very beginning of the simulation because the inflow does not have to overcome an angular momentum barrier as it does not possess a significant rotational velocity. For $T\gtrsim 650\,r_{\mathrm{g}}/c$, we find that both $\Phi_{\mathrm{B}}$ and $\dot{M}$ start to saturate.
@@
Figure 3: Growth of a Kelvin-Helmholtz-like mode in the jet-disk boundary over time, where effective mixing between the disk ions ($i$) and the jet positrons ($e^{+}$) occurs as shown by the number density fraction ($n_{e^{+}}/n_{i}$). The swirl-like motion of the KH vortices is clearly denoted by the three-velocity streamlines. The Z coordinates vary for the different panels and are defined by $\mathrm{Z}_{\mathrm{cent}}+\Delta\mathrm{Z}$. The shown instances are time-averaged over a window of $\Delta\mathrm{T}\approx 7\,r_{\mathrm{g}}/c$.
@@
Fig. 2 shows the characteristics of the MAD accretion state in our collisionless simulation. Panel 2a shows a still on-going flux eruption where the accretion flow is receding and the magnetic field lines are reconnecting. Panel 2b displays the moment when the accretion flow reinstates itself, creating a clear kink mode in the equatorial current sheet [a manifestation of the relativistic drift-kink; 53, 54]. The disk-magnetosphere system can be divided into several core regions that are of dynamical interest: (i) the equatorial current sheet, (ii) the jet sheath (i.e., the jet-disk interface), (iii) the inner (BZ) jet, (iv) the polar inner magnetosphere (enclosing the spark gap), and (v) the disk. The overdense equatorial plane is the result of the enclosing properties of the poloidal magnetic-field loop structure. Moreover, the tearing modes of the equatorial current sheet [28, 29, 55] instigate the creation of pairs, as it coincides with particle acceleration due to non-ideal electric fields at X-points associated with magnetic reconnection, which contributes to the overdense regions in the equatorial plane.
@@
We find that the flux eruptions occasionally result in the stretching and curling of magnetic fields along the jet sheath. This is caused by an initially non- or slowly rotating magnetic field line that is suddenly set into motion after it enters the BH ergosphere. The sudden jolt of the magnetic field line travels outwards along the jet sheath and creates sufficient velocity shear to initiate KH-like behavior. Some of these wave roll-ups [see also 30] eventually form circular (island) magnetic field structures [57, 55]. These structures grow via a swirl-like flow which drags magnetic fields that relax into a circular shell around the original structure.
@@
Fig. 1d shows the evolution of $\gamma_{\mathrm{max}}$ over time (which is also spatially identifiable with the $\langle\xi\rangle$ variable in Fig. 2), where it becomes clear that $\gamma_{\mathrm{max},e^{+}}\gtrsim\gamma_{\mathrm{max},e^{-}}$ during selected flux eruption events (i.e., the vertical lines). These events coincide with the restoration of the accretion flow after a flux eruption. This connects to the findings outlined in Fig. 4 [and Supplemental Material; 48], where we establish the spatial locations of $\gamma_{\mathrm{max}}$ per particle species. These numbers should be evaluated as a fraction of the limiting Lorentz factor implied by the potential drop $\hat{\gamma}_{\mathrm{max}}=ea_{\ast}Br_{\mathrm{g}}/m_{e}c^{2}\simeq 2\times 10^{5}$ [32] taking into account the increase in $B$ over time. This is verified in Fig. 1b where $\Phi_{\mathrm{B}}$ increases ${\sim}20$ times with respect to the starting value, pushing the simulation closer to a realistic separation of scales. The inner polar magnetosphere regions produces the highest $\gamma_{\mathrm{max}}\sim 10^{5}$ for the electrons and, to a lesser degree, also for the positrons. The ions do not exceed $\gamma_{\mathrm{max},i}\sim 10^{3}$, but we note they do reach the same total energy when one compensates for the mass difference as $m_{i}\gamma_{\mathrm{max},i}/m_{e}\sim 10^{5}$. We note that $\gamma_{\mathrm{max},e^{-}}$ and $\gamma_{\mathrm{max},e^{+}}$ can originate in different regions, where the positrons are mainly accelerated in the equatorial reconnection region. The timing of this acceleration coincides with the restoration of the accretion flow after a reconnection event. The pairs created at the X-point are collected at the front of the restoring accretion flow. Occasionally, a considerable (vacuum) spark gap forms on a similar time-scale, which can then affect and slightly oppose the restoration of the accretion flow (see Fig. 2b).
@@
The most accelerated electrons originate within the inner magnetosphere region where pair creation dynamics associated with the spark gap are known to be important [33]. At the high optical depth to pair creation used here, gap particles are created in intermittent bursts of pairs originating close to the BH that are then accelerated outward (i.e., electrons) or inward (i.e., positrons). Fig. 2 shows the two distinct channels for the pair production cascade, which can be advected into the equatorial region (mainly accelerating positrons, panel 2b) or along the polar (jet) regions (mainly accelerating electrons, panel 2a). Especially the equatorial pair cascade (associated with significant unscreened electric fields [48]), which creates a bubble, has not been seen before. For the modest multiplicity pair plasma (i.e., $n_{e^{\pm}}/n_{\mathrm{GJ}}\lesssim 5$), we note that the roles (i.e., inflowing/outflowing) of the leptons are primarily determined by the choice in the sign of $\bm{\Omega}_{\mathrm{h}}\cdot\bm{B}$; at higher multiplicities this asymmetry should vanish, and polar-cap acceleration should be highly diminished while reconnection-driven particle acceleration should remain mostly unchanged.
@@
Figure 4: Particle energy distribution from distinct accretion flow regions for the ion ($i$, solid), electron ($e^{-}$, dashed), and positrons ($e^{+}$, dotted lines). We show the decomposed spectra associated with the two (out of five) regions, which correspond to equatorial reconnection layer (rec) and inner polar BH magnetosphere (inn). These are the regions where the particles are most accelerated. The spectrum is averaged over the time window $T\in[726.85,736.14]\,r_{\mathrm{g}}/c$. The vertical lines denote the $\gamma_{\mathrm{max}}$ values that disregard the tail of the distribution. Details of the underlying procedures are described in the Supplemental Material [48].

### v2

@@ Abstract
We present the first collisionless realization of two-dimensional axisymmetric black hole accretion consistent with a persistent magnetically arrested disk state. The accretion flow, consisting of an ion-electron disk plasma combined with magnetospheric pair creation effects, is simulated using first-principles general-relativistic particle-in-cell methods. The simulation is evolved over significant dynamical timescales during which a quasi-steady accretion state is reached with several magnetic flux eruption cycles. We include a realistic treatment of inverse Compton scattering and pair production, which allows for studying the interaction between the collisionless accretion flow and pair-loaded jet. Our findings indicate that magnetic flux eruptions associated with equatorial magnetic reconnection within the black hole magnetosphere and the formation of spark gaps are locations of maximal particle acceleration. Flux eruptions, starting near the central black hole, can trigger Kelvin-Helmholtz-like vortices at the jet-disk interface that facilitate efficient mixing between disk and jet plasma in this region. Transient periods of increased pair production following magnetic flux eruptions and reconnection events are responsible for most of the highly accelerated particles.
@@
Supermassive black holes (BH) at the center of low-luminosity active galactic nuclei (LLAGN) accrete ion-lepton plasma whose high temperatures and low densities place it in the collisionless regime  [^1] The collisionless regime refers to the mean free path to Coulomb collisions being large – often exceeding the system size by a multitude of factors [e.g, 65]. Some of these BHs are known to launch jets, but these features are well-established to not be universal [2, 3, 4]. Examples can be found in the two LLAGN M87 and Sagittarius A, where M87 shows a prominent, large-scale jet and Sagittarius A only displays radiative characteristics typical for a compact jet [5, 6]. The Event Horizon Telescope has observed emission structures associated with millimeter-wavelength synchrotron radiation from near-horizon relativistic accretion flows for both systems [7, 8]. Although these emission features are likely produced by the bulk accretion flow, the question remains of how and where the higher-energy emission and particles are created.
@@
Many jetted black hole systems, including M87 and Sagittarius A, display $\gamma$-ray activity [9, 10, 11, 12, 13, 14, 15]. These systems can go through periods of increased (i.e., flaring) activity [16, 17, 18], which can be related to the dynamics of the central engine that drives jet launching mechanics and its associated plasma effects. The source of the high-energy emission is often tied to the collisionless plasma dynamics within the BH magnetosphere, but it remains a challenge to identify the responsible mechanisms and locations. With this numerical study, we therefore aim to simultaneously assess the ability of accretion-driven and magnetospheric processes to accelerate particles to high, non-thermal energies.
@@
The Magnetically Arrested Disk [MAD; 19, 20] state is currently a preferred model for interpreting millimeter emission from the two aforementioned LLAGN, both with regard to the polarimetric signatures [21] and the origin of flares with associated orbital motion of hot spots [22, 23, 24, 25]. MAD episodes are characterized by flux eruptions, during which a limiting horizon-penetrating magnetic flux is reached, or even exceeded, resulting in the partial halting of the accretion flow [e.g., 26, 27, 28, 29, 30, 31]. General-relativistic magnetohydrodynamical (GRMHD) models that bring about the MAD state are characterized by large-scale coherent magnetic fields and sizable initialized disk structures. However, GRMHD methods are technically not applicable in the collisionless regime and do not directly provide insight into the non-ideal processes associated with particle acceleration. It is therefore valuable to investigate these systems with methods that self-consistently take into account these processes while simultaneously building on the GRMHD findings.
@@
In this Letter, we employ general relativistic (radiative) particle-in-cell (GR[R]PIC) methods to undertake a global collisionless study combining initialized accretion flows and particle creation processes. We present a quasi-steady, axisymmetric MAD accretion state and focus on interpreting the locations and mechanisms of particle acceleration within this environment. The system’s characteristics — highly magnetized, with large-scale coherent field structures — are reminiscent of the well-established MAD state in GRMHD models [e.g., 26, 27, 28, 29, 30, 31]. Previously, several kinetic studies of black hole magnetospheres were undertaken that mostly focused on pair-created plasma dynamics under magnetically dominated conditions [32, 33, 34, 35, 36, 37, 38] or spherically-symmetric Bondi-Hoyle accretion [39]. Here, we present the first realization of disk-like initial conditions in a collisionless scenario, consisting of a three-species (pair-ion) plasma with a significant ion-to-electron mass ratio. With the addition of a strong accretion flow component to the BH magnetosphere in the presented simulation, it becomes possible to better constrain the flaring dynamics of AGN systems and gain insight into the edge-brightened mechanisms of AGN jets [40, 41, 42, 43, 44, 45].
@@
We use the GRRPIC code Zeltron [46, 32] to evolve an ion-electron accretion flow combined with a pair-filled black hole magnetosphere. The total simulation domain is denoted by $r\in[0.9r_{\mathrm{h}},27r_{\mathrm{g}}]$ and $\theta\in[0.01\pi,0.99\pi]$ with a resolution of $N_{r}\times N_{\theta}=1440^{2}$ for spherical Kerr-Schild coordinates with a grid that is logarithmically spaced in $\hat{r}$ and linearly spaced in $\hat{\theta}$. Here, $r_{\mathrm{g}}=GM/c^{2}$ is the gravitational radius and $r_{\mathrm{h}}=(1+\sqrt{1-a^{2}})r_{\mathrm{g}}$ is the horizon radius, with gravitational constant $G$, black hole mass $M$, speed of light $c$, and dimensionless black hole spin $a=0.99$. The ion mass is set to $m_{i}=100\,m_{e}$, where $m_{e}$ is the electron mass. We initialize a uniform number density, $n=8\,n_{\textrm{GJ}}$, wedge that spans $r\in[5,27]r_{\mathrm{g}}$, $\theta\in[\pi/3,2\pi/3]$, where $n_{\mathrm{GJ}}=B_{0}\omega_{\mathrm{h}}/(4\pi ce)$ is the Goldreich-Julian density with angular velocity $\omega_{\mathrm{h}}=ac/(2r_{\mathrm{h}})$ [47] and $B_{0}$ the initial magnetic field strength [see Sup. Mat.; 48]. In the inner region, i.e., $r\in[r_{\mathrm{h}},5\,r_{\mathrm{g}}]$, we initialize a uniform pair-plasma sphere with a number density of $1\,n_{\mathrm{GJ}}$ to kickstart pair production near the black hole. The magnetic field begins as a single large magnetic loop in the poloidal plane [see Sup. Mat.; 48] with an initialized ratio of thermal to magnetic pressure $\beta\in[0.1,5]$. Our primary focus is investigating the collisionless implications of a persistent MAD accretion state.
@@
The initial configuration of the ion-electron plasma is somewhat arbitrarily chosen and of little consequence to the formation of the relaxed, quasi-steady axisymmetric state that is achieved after the simulation has evolved for sufficient time [see Sup. Mat.; 48]. No angular momentum is introduced for the accretion flow at initialization, which expedites the onset of an axisymmetric quasi-steady accretion state. However, we note that frame-dragging effects will quickly introduce angular momentum for the innermost black hole regions. For simplicity, we refer to the low angular-momentum accretion flow as the “disk”. The ion-electron disk plasma is initialized with a temperature $k_{\mathrm{B}}T_{\mathrm{inj}}=m_{e}c^{2}$ for both ions and electrons. Plasma is replenished from a spherical shell corresponding to $r\in[25,27]r_{\mathrm{g}}$, $\theta\in[\pi/3,2\pi/3]$, where injection occurs at a fixed rate when the density drops below the initialized disk density. The injection from the boundary emulates the inflow of matter from a large-scale accretion flow. Particles that pass through the outer boundary are removed [49].
@@
Electron-positron pairs are created via photon-photon annihilation of inverse-Compton scattered $\gamma$-ray photons with a background photon field [48]. The inclusion of these particle creation effects provides insight into how and if the jet is loaded with pairs and whether these pairs are able to significantly interact at the jet-disk interface. We refer the reader to the Supplemental Material [48] (and references [50, 33]) for a more detailed description of the simulation specifics, such as the definition of the poloidal magnetic field initialization, pair production criteria, and characteristic scaling.
@@
Figure 2: Overview of the axisymmetric MAD accretion state during and after a flux eruption. The displayed quantities are magnetization $\sigma_{\mathrm{h}}=B^{2}/\left[4\pi\sum_{s}\langle\gamma_{s}\rangle n_{s}m_{s}c^{2}\right]$, where $\langle\gamma_{s}\rangle$ is the cell-averaged Lorentz factor and $\rho_{s}=m_{s}\,n_{s}$ the mass density per species, $s\in\{i,e^{-},e^{+}\}$, normalized with the Goldreich-Julian mass density $\rho_{\mathrm{GJ}}=m_{e}\,n_{\mathrm{GJ}}$. Kelvin-Helmholtz vortices are seen to form along the jet-disk interface. The inset panels show the density-averaged energy per particle $\langle\xi\rangle=\sum_{s}\langle\gamma_{s}\rangle\rho_{s}c^{2}/\sum_{s}\rho_{s}c^{2}$ for different regions of the simulation domain. The green contours outline the magnetic flux surfaces. The dashed blue contours denote the ergosphere boundary.
@@
The accretion flow dynamic is well summarized in the time series of horizon-penetrating fluxes, where we define the mass accretion rate $\dot{M}=-\iint_{A}\rho u^{r}\sqrt{-g}\mathrm{d}\theta\mathrm{d}\phi$ and magnetic flux $\Phi_{\mathrm{B}}=\tfrac{1}{2}\iint_{A}|B^{r}|\sqrt{h}\mathrm{d}\theta\mathrm{d}\phi$ on the horizon ($r=r_{\mathrm{h}}$). Here, $\rho$, $u^{\mu}$, $g$, $h$, and $B^{r}$ are the mass density, four-velocity, metric determinant, spatial metric determinant, and the radial component of magnetic field as seen by local fiducial observers (FIDOs) that are normal to spatial hypersurfaces, respectively. The quantity known as the MAD parameter, defined as $\phi_{\mathrm{BH}}=\Phi_{\mathrm{B}}/\sqrt{\dot{M}r^{2}_{\mathrm{g}}c}$, is known to saturate at ${\sim}50$ [51]. After it saturates, a flux eruption event occurs, where the magnetic flux is (partially) pushed away from the BH. This leads to reconnection of the upper and lower hemispheres of the BH magnetosphere in the equatorial plane [29, 30, 31], and a resulting drop in $\Phi_{\mathrm{B}}$. We also find that the pairs created by photon-photon annihilation leave an imprint on the system’s flux eruptions, as a significant (vacuum) gap can form during the peak of the flux eruption that is rapidly filled by a pair production cascade (Fig. 2b).
@@
In Fig. 1, we show direct evidence for a quasi-steady axisymmetric MAD accretion state achieved with GRRPIC methods. The top panel clearly outlines sharp decreases in the mass accretion rate $\dot{M}$, which are direct consequences of the flux eruption events, where the accretion flow is temporarily halted because of a magnetic field barrier created by the flux eruption. This relates to the sawtooth structure seen for $\Phi_{\mathrm{B}}$, where the exponential decay is related to the reconnection event (or flux eruption). The general shape of growing $\Phi_{\mathrm{B}}$ with the sawtooth structures of ${\sim}50\,r_{\mathrm{g}}/c$ in length combined with significant amplitude changes have not been seen before in other GRPIC simulations [cf. 34, 52, 39]. The MAD state, corresponding to reaching the limit $\phi_{\mathrm{BH}}\gtrsim 50$, is established from the very beginning of the simulation. For $T\gtrsim 650\,r_{\mathrm{g}}/c$, we find that both $\Phi_{\mathrm{B}}$ and $\dot{M}$ start to saturate.
@@
Figure 3: Growth of a Kelvin-Helmholtz-like mode in the jet-disk boundary over time, where mixing between the disk ions ($i$) and the jet positrons ($e^{+}$) occurs as shown by the number density fraction ($n_{e^{+}}/n_{i}$). The swirl-like motion of the KH vortices is clearly denoted by the three-velocity streamlines. The streamline thickness is scaled according to the magnitude of the velocity. The panels are shown as a function of $\mathrm{Z}^{\prime}=\mathrm{Z}-\mathrm{Z}_{\mathrm{cent}}$. The shown instances are time-averaged over a window of $\Delta\mathrm{T}\approx 7\,r_{\mathrm{g}}/c$.
@@
Fig. 2 shows the characteristics of the MAD accretion state in our collisionless simulation. Panel 2a shows a still ongoing flux eruption where the accretion flow is receding and the magnetic field lines are reconnecting. Panel 2b displays the moment when the accretion flow reinstates itself, creating a clear kink in the equatorial current sheet [a manifestation of the relativistic drift-kink instability; 53, 54]. The disk-magnetosphere system can be divided into several core regions that are of dynamical interest: (i) the inner jet, (ii) the jet sheath (i.e., the jet-disk interface), (iii) the equatorial reconnection region, (iv) the inner (polar) magnetosphere (enclosing the spark gap), and (v) the disk. The overdense equatorial plane is in part the result of the enclosing properties of the poloidal magnetic-field loop structure. Moreover, the tearing modes of the equatorial current sheet instigate the creation of pairs, which further contributes to the overdense equatorial plane region. These pairs are created in the vicinity of X-points (locations of significant non-ideal electric fields) of the magnetically reconnecting layer, where efficient particle acceleration is known to occur [28, 29, 55].
@@
We find that the flux eruptions occasionally result in the stretching and curling of magnetic fields along the jet sheath. This is caused by an initially non- or slowly rotating magnetic field line that is suddenly set into motion after it enters the BH ergosphere creating a torsional Alfvén wave. The oscillations of the Alfvén wave result in sufficient velocity shear to bring about KH-like behavior. Some of the KH eddies [see also 30] eventually form circular (island) magnetic field structures [57, 55]. These structures then grow via the surrounding swirl-like flows which drag magnetic fields that relax into a circular shell around the original structure.
@@
Fig. 1d shows the evolution of $\gamma_{\mathrm{max}}$ over time (which is also spatially identifiable with $\langle\xi\rangle$ in Fig. 2), where it becomes clear that $\gamma_{\mathrm{max},e^{+}}\gtrsim\gamma_{\mathrm{max},e^{-}}$ during selected flux eruption events (i.e., the vertical lines). These events coincide with the restoration of the accretion flow after a flux eruption. This connects to the findings outlined in Fig. 4 [and Sup. Mat.; 48], where we establish the spatial locations of $\gamma_{\mathrm{max}}$ per particle species. These numbers should be evaluated as a fraction of the maximal Lorentz factor implied by the potential drop $\hat{\gamma}_{\mathrm{max}}=eaBr_{\mathrm{g}}/m_{e}c^{2}\simeq 2\times 10^{5}$ [32], taking into account the increase in $B$ over time. This is verified in Fig. 1b where $\Phi_{\mathrm{B}}$ increases ${\sim}20$ times with respect to the starting value, pushing the simulation closer to a realistic separation of scales. The inner polar magnetosphere regions produces the highest $\gamma_{\mathrm{max}}\sim 10^{5}$ for the electrons and, to a lesser degree, also for the positrons. The ions plateau around $\gamma_{\mathrm{max},i}\sim 10^{3}$, but they do reach the same total energy when one compensates for the mass difference as $m_{i}\gamma_{\mathrm{max},i}/m_{e}\sim 10^{5}$. We note that $\gamma_{\mathrm{max},e^{-}}$ and $\gamma_{\mathrm{max},e^{+}}$ can originate in different regions, where the positrons are mainly accelerated in the equatorial reconnection region. The timing of this acceleration coincides with the restoration of the accretion flow after a reconnection event. The pairs created near the X-point are collected at the front of the restoring accretion flow. Occasionally, a considerable (vacuum) gap forms on a similar time-scale, which is rapidly filled with pairs that can then affect and slightly oppose the restoration of the accretion flow (Fig. 2b).
@@
The most accelerated electrons originate within and near the ergosphere region where pair creation dynamics associated with spark gaps are known to be important [33]. At the high optical depth to pair creation used here, the gap is small (${\sim}0.01r_{\mathrm{g}}$) and pairs are created in intermittent bursts originating close to the BH that are then accelerated outward (i.e., electrons) or inward (i.e., positrons). Fig. 2 shows the two distinct channels for the pair production cascade, which can be advected into the equatorial region (mainly accelerating positrons, Fig. 2b) or along the polar (jet) regions (mainly accelerating electrons, Fig. 2a). Especially the equatorial pair cascade (associated with significant unscreened electric fields [48]), creating an equatorial evacuated (bubble-shaped) structure (Fig. 2b), has not been seen before. For the modest multiplicity pair plasma (i.e., $n_{e^{\pm}}/n_{\mathrm{GJ}}\lesssim 5$), we note that the roles (i.e., inflowing/outflowing) of the leptons are primarily determined by the choice in the sign of $\bm{\Omega}_{\mathrm{h}}\cdot\bm{B}$; at higher multiplicities, this asymmetry should vanish and polar-cap acceleration should be highly diminished while reconnection-driven particle acceleration should remain mostly unchanged.
@@
Figure 4: Two most prominent particle acceleration regions correspond to the equatorial reconnection layer (rec) and inner polar BH magnetosphere (inn). Particle energy distribution from distinct accretion flow regions for the ion ($i$, solid), electron ($e^{-}$, dashed), and positrons ($e^{+}$, dotted lines). The thick lines correspond to the spectra calculated for all particles in the $r\leq 25r_{\mathrm{g}}$ domain. The spectra are averaged over the time window $T\in[726.85,736.14]\,r_{\mathrm{g}}/c$ corresponding to Fig. 2a. The vertical lines denote the $\gamma_{\mathrm{max}}$ values that disregard the tail of the distribution. More details of the spectra are outlined in the Sup. Mat. [48].